A car pulls out from a stop sign and builds up to a speed of 20.0m/sec in 8 seconds. How far does it travel in that time? What final speed does it reach?
Answer by Alan3354(69443)

A car pulls out from a stop sign and builds up to a speed of 20.0m/sec in 8 seconds. What final speed does it reach?
It's stated that it's 20 m/sec ??
a = 20/8 = 2.5 m/sec/sec acceleration
--------------------
How far does it travel in that time?
Its average speed = (0 + 20)/2 = 10 m/sec
10*8 = 80 meters